FIS2005 http://www.mdpi.org/fis2005/

This is the title of my paper.

Michel Petitjean ITODYS (CNRS-Université Paris 7, UMR 7086), 1 rue Guy de la Brosse, 75005 Paris, France E-mail: petitjean@itodys.jussieu.fr URL: http://petitjeanmichel.free.fr/itoweb.petitjean.html

Abstract: Here is the abstract of my paper for the Third Conference on the Foundations of Information Science. Keywords: data mining; knowledge data discovery; MSC 2000 codes: optional optional PACS codes: optional optional

\begin{matrix}c \\ ◯ \end{matrix}2005 by MDPI (http://www.mdpi.org). Reproduction for noncommercial purposes permitted.

1 Introduction

Introduction...

2 Method

Theory...

3 Results

Results...

4 Conclusion

Conclusion...

5 References

[1] ...